A 6-in bare oxidized steel pipe (6.625 in O.D) carries steam at a temperature of 540°F and is loacted in a large room having wall temperatures of 80 °F. Determine the net radiant-heat exchange in BTU/(hr)(ft) between the pipe and the room, (a) if the surface temperature of the pipe is assumed to be the same as the steam temperature; (b) if the pipe is covered with a single layer of asbestos paper and the temperature of the outer surface of the paper is 530 °F; (c) if the asbestos paper painted with aluminum paint containing 26 percent aluminum and the surface temperature is 530 °F.
1
Expert's answer
2021-10-29T02:33:08-0400

a)

"T_1=540"

"T_2=80"

"\epsilon_1=0.9"

"\epsilon_2=0.4"

"\sigma=5.67*10^{-8}"

"Q_{net}=\frac{\sigma(T_1^4-T_2^4)}{\frac{1}{\epsilon_1}+\frac{1}{\epsilon_2}-1}"

"Q_{net}=\frac{5.67*10^{-8}(540^4-80^4)}{\frac{1}{0.9}+\frac{1}{0.4}-1}"

"=1845.54 BTU\/hr ft"


b)

"T_1=530"

"T_2=80"

"\epsilon_1=0.9"

"\epsilon_2=0.4"

"\sigma=5.67*10^{-8}"

"Q_{net}=\frac{\sigma(T_1^4-T_2^4)}{\frac{1}{\epsilon_1}+\frac{1}{\epsilon_2}-1}"

"Q_{net}=\frac{5.67*10^{-8}(530^4-80^4)}{\frac{1}{0.9}+\frac{1}{0.4}-1}"

="1712.52" BTU/hr ft


c)"T_1=530"

"T_2=80"

"\epsilon_1=0.67"

"\epsilon_2=0.4"

"\sigma=5.67*10^{-8}"

"Q_{net}=\frac{\sigma(T_1^4-T_2^4)}{\frac{1}{\epsilon_1}+\frac{1}{\epsilon_2}-1}"

"Q_{net}=\frac{5.67*10^{-8}(530^4-80^4)}{\frac{1}{0.67}+\frac{1}{0.4}-1}"

"=1494.24" BTU/hr ft